A ball is projected with a speed of 20m/sec at an angle of 30 degree from a point on top of very high tower. The time after which it`s velocity becomes perpendicular to the the velocity of projection is

Grade:9

1 Answers

Viresh kumar
6 years ago
Solution to the question is very easy,just think logically...
Particle is thrown at angle 30 with horizontal...
AT the time when it’s velocity is perpendicular to initial velocity,its velocity will be making an angle 120 with the positive x-axis..
THen,by using v=u+at,
20.sin120=20.sin30 – 10t (a=-g = -10)
-10.31/2 = 10 – 10t
t= 1 + 31/2
t= 2.73 s